Grade 12th passBranch Ranking of IIT

if i opt to take ece as my branch..then which iit campus should i choose which is giving me this opportunity?
My expected score is 213/366 for jee advanced

Rukalpa De
9 Years agoGrade 12th pass

1 Answer

Vasanth SR
8 Years ago
RANK = 10001-20000 Marks = 190-218
With this IIT Kanpur and IIT Roorkee are possible